Joshua is one-third as old as his father who is three years older than his mother. When Joshua was born the sum of the ages of his father and mother was 53. How old is Joshua now? ( answer is 14 - show your solution )
Answer by mananth(16946)

Joshua is one-third as old as his father who is three years older than his mother. When Joshua was born the sum of the ages of his father and mother was 53. How old is Joshua now? ( answer is 14 - show your solution )
mother's age be x
father's age = x+3
Joshua's age = (x+3)/3
Father's age when Joshua was born (x+3)- (x+3)/3 = (2x+6)/3
Mother's age when Joshua was born x-(x+3)/3= (2x-3)/3

(2x+6+2x-3)/3 = 53
4x+3 = 159
4x= 159-3
4x=156
x=39 mother's age
fathers age = 42
Joshua's age = 42 /3 = 14 years